OpenStudy (anonymous):

URGENT!!!: What is the range of f^-1[g(x)]???

sammixboo (sammixboo):

Maybe @triciaal or @jim_thompson5910 could help :)

jimthompson5910 (jim_thompson5910):

There seems to be missing information. Can you post a screenshot?

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

here!!

jimthompson5910 (jim_thompson5910):

so this is all part of #6 and \[\Large f(x) = \frac{2}{x+3}\] \[\Large g(x) = 2^{x-1}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im stuck on part b). i have the domain i just cant figure out the range

jimthompson5910 (jim_thompson5910):

I agree that \[\Large f^{-1}(x) = \frac{2-3x}{x}\] \[\Large f^{-1}(g(x)) = \frac{2-3(2^{x-1})}{2^{x-1}}\]

jimthompson5910 (jim_thompson5910):

you have the correct domain as well

jimthompson5910 (jim_thompson5910):

as for the range, let me think

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

\[\Large f^{-1}(x) = \frac{2-3x}{x}\] \[\Large f^{-1}(x) = \frac{2}{x} - \frac{3x}{x}\] \[\Large f^{-1}(x) = \frac{2}{x} - 3\] \[\Large f^{-1}(g(x)) = \frac{2}{2^{x-1}} - 3\]

jimthompson5910 (jim_thompson5910):

as x gets larger and larger, the term \(\LARGE \frac{2}{2^{x-1}}\) gets smaller and smaller effectively getting to small that it becomes 0

jimthompson5910 (jim_thompson5910):

so as x gets larger and larger \[\Large f^{-1}(g(x)) = \frac{2}{2^{x-1}} - 3\] gets closer and closer to -3. It will never actually get to -3 itself though

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (anonymous):

wait so whats the range?

jimthompson5910 (jim_thompson5910):

if you plug in x = 0, what do you get?

OpenStudy (anonymous):

idont know :(((((

OpenStudy (anonymous):

@triciaal

jimthompson5910 (jim_thompson5910):

\[\Large f^{-1}(g(x)) = \frac{2}{2^{x-1}} - 3\] \[\Large f^{-1}(g(0)) = \frac{2}{2^{0-1}} - 3\] \[\Large f^{-1}(g(0)) = ???\]

OpenStudy (triciaal):

consider when x is negative when x = 0 and when x is positive

OpenStudy (triciaal):

you said you had the actual domain the range is the set of y values

OpenStudy (anonymous):

can i pls just have the answer pls. this homework q is the last one and im rly tired plz

OpenStudy (triciaal):

maybe -infinity to -3 not sure

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

if you graph the entire thing, you'll get this

jimthompson5910 (jim_thompson5910):

it goes up forever but doesn't do the same in the other direction. Instead it gets closer and closer to -3 (but never actually gets to -3) so the range is y > -3 which in interval notation is (-3, infinity)

OpenStudy (anonymous):

omg thank you guys so much

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

i can finally sleep!!!!!

OpenStudy (triciaal):

I see my error when x = -infinity y is positive so the solution is like Jim has it

OpenStudy (anonymous):

guys!!!

OpenStudy (anonymous):

is anyone here still able to help me with one more q?

jimthompson5910 (jim_thompson5910):

go ahead